Summary of the Meeting of the PS
held on 13.05.2002
Present : B. Autin, S. Baird, J.
Boillot, M. Chanel, B. Frammery, R. Garoby, G. Geschonke, J. Gruber,
C. Hill, T. Kehrer (Secretary), E. Kugler, K. Metzmacher, J.P.
Potier, J.P. Riunaud, K. Schindl (Chairman), I. Wilson
Excused : B.W. Allardyce, A. Blas (Staff Association), G. Daems, J.P.
Delahaye, H. Haseroth, M. Lindroos, S. Maury, S.L. Neboux
* * * * *
1. Follow-up of previous meeting and matters arising
- Restructuring of the Accelerator Sector
- The name proposed for the new PS/SL Division is "Accelerators
& Beams" (AB).
- The Dumps & Targets group will stay in the AB Division and will
not be moved to EST (Technical Sector) as earlier proposed.
- J.-P. Delahaye suggests the Group Leaders to announce to their groups
- the structure of the new AB Division;
- the proposed Group Leaders (these nominations will only become
official after the Council of June 2002 and should not be mentioned
in any document until that date);
- That the new A&B division will take the best advantage of
the skill and qualities from both PS and SL staff, with a fair and
equal distribution of responsibilities and without any division
taking over the other one;
- Removals will be strictly limited.
- PS/SL Management Meeting, 27-29 May : the three-day meeting
will take place in Morzine. Arrival on Sunday evening or Monday morning.
The "old" and "new" management, the DPOs and the
DAOs of both divisions will attend. The programme is not yet defined
and expected by the end of this week. There is no transport organized
so people are requested to take their own cars.
- Restructuring of Controls : the SLTC (SL Technical Committee)
organizes 2 meetings (Wednesday and Thursday 15 and 16 May 2002 - Agenda) on the structure
and task of the future common Control Group.
- Linac flooding (Action S. Baird) : whilst ST Division is looking
for a permanent solution, the leaky heat exchanger in the Linac air conditioning
system was temporarily repaired.
- Meetings : due to Whitsun on Monday 20 May 2002, the next Schedule
Meeting will take place Tuesday 21 May 2002 at 10:00 and the PSMB meeting
Tuesday 21 May 2002 at 14:30.
- Ions for LHC External Review : this review scheduled for August
will be postponed to September 2002.
- Safety : C. Hill reported in the absence of G. Daems on safety
- Asbestos in South Hall (LEIR Ring) : An external specialist
from the Geneva cantonal office (Ecotox) has been consulted. The results
of asbestos fibre concentration analysis are slightly higher that the
Federal limit but as the precision of this type of measurement is rather
low near this limit, further tests will be done to determine if it is
a matter of local or general contamination.
- Power supply of septum 57 : following the incident of 2nd May
2002 on this high voltage power supply, an accident report has been
published and sent to all parties concerned. While no particular short-term
action is expected from the Division the situation and Staff awareness
has to be markedly improved in the medium term. The Safety Instructions
must be followed at all times.
- Pipe-work in the PS : while the asbestos is now removed from
the pipes which have been taken out from the tunnel, there remains the
issue of how to dispose of this slightly radioactive material. TIS would
like to get PS Division involved, but the project is entirely under
ST responsibility. The DSO will clarify the situation (action : G.
- LHC beam for SPS beam scrubbing tests : M. Chanel explains why
transverse emittances are too big (3.5 mm at
180 1010 instead of 2.5 mm 130 1010
proton per PSB ring). R. Garoby informs the meeting that the setting-up
is not finished yet and that there are still problems (ex. cavity problem
over the weekend, excessive losses between PSB and SPS) and that due to
the other required beams (AD, nTOF) people work under very high pressure.
S. Baird recalled that the PS gave the first beam to the SPS already on
Wednesday 8 May 2002 instead of the originally scheduled date of today.
- AD : the AD startup was
initially scheduled on Thursday 23 May 2002, but will most probably
be postponed until Monday 27 May 2002 due to overcharge of the PS
- ADUC : next meeting will take place Friday afternoon 17 May 2002.
- ISOLDE : runs fine, 2
separators (GPS, HRS) in parallel.
- CTF3 : RF recombination
kickers will be installed this Friday 17 May 2002. The first tests
are planned next week, after Whitsun.
Division | PSMB Meetings | Other PS Meetings |
tk 13 May 2002